OpenStudy (anonymous):

In 2003, Company A reported sales of $2 million. In 2004, the company had sales of $1.2 million. Find the percent of change to the nearest tenth of a percent. Is this a percent of increase or decrease?

OpenStudy (swissgirl):

omg what is this for? All ur questions r soooo different

OpenStudy (anonymous):

@swissgirl its a bunch of things mixed together...do u know the answr?

OpenStudy (swissgirl):

Like idk I am getting a bad number

OpenStudy (swissgirl):

I was getting 66.66666667 %

OpenStudy (anonymous):

thanks!!

OpenStudy (swissgirl):

I am not vouching for this

OpenStudy (kinggeorge):

I think you need to calculate the number \[\frac{1.2-2}{2}\]to get your percentage. Just looking at it, this should be less than 50%.

OpenStudy (swissgirl):

http://www.mathsisfun.com/numbers/percentage-change.html I just followed this but ur seems like a better option

OpenStudy (kinggeorge):

How did you end up getting 66.7%?

OpenStudy (swissgirl):

Step 1: Divide the New Value by the Old Value (you will get a decimal number) Step 2: Convert that to a percentage (by multiplying by 100 and adding a "%" sign) Step 3: Subtract 100% from that so 2/1.2*100-100

OpenStudy (kinggeorge):

That should give you the same thing that I had.

OpenStudy (anonymous):

(2-1.2)/2 * 100%

OpenStudy (swissgirl):

Panlac Ik I am wroonngggg

OpenStudy (kinggeorge):

It needs to be \((1.2-2)\) in the numerator since you want new-old.

OpenStudy (anonymous):

opps typed it wrong

OpenStudy (anonymous):

1.2 - 2

OpenStudy (anonymous):

any value that depreciated must have a negative percentage

OpenStudy (swissgirl):

@sushi323 is long gone and on to her next question i assume lol

OpenStudy (anonymous):

it's the same movement in slopes

OpenStudy (swissgirl):

trueee

OpenStudy (anonymous):

now clean it up and let us move on

OpenStudy (anonymous):

hahaha yepp

OpenStudy (kinggeorge):

the % sign begins a comment in Latex, so everything after it doesn't appear.

OpenStudy (anonymous):

(1.2 - 2)/2 * 100% = x

OpenStudy (anonymous):

thanks everyone i really appreciate it!!!

OpenStudy (kinggeorge):

\[\frac{1.2 - 2}{2} \cdot 100 \% = x\]There we go.